Ticket Number: HEA-NOC/20041004-1 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver: TCD
Ticket Opened: 20041004 07:25 UTC+1 Problem Start: 20041003 21:05 UTC+1
Ticket Update: 20041006 09:59 UTC+1 Problem End:
Hits on TCD primary access, Colt GE LE-001892.
Loss of connectivity with TCD.
20041004 07:19 MN Check BH access router, logs show outage from 21:05 to
23:30, and now from 06:34. Report to Colt, ticket #448068 issued.
20041004 07:54 MN Line comes up at 07:48. Check site router: it shows
cycle at 07:48. Advise Colt.
20041004 11:11 BN Line has been up since 07:48, line appears to be running
clean. Crashdump info has been passed on to Lancomms and Cisco TAC are
investigating. Connection should still be considered at risk.
20041006 09:57 BN Router instability occured this morning during
maintenance window. Connection was eventually restored and crash
information is being gathered.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20041004-1